Ron and Kathy are ticket-sellers at their class play, Ron handling student tickets that sell for $2.00 each and Kathy selling adult tickets for $4.50 each. If their total income for 20 tickets was $60.00, how many did Ron sell?
Answer by Maths68(1474)

Let
Student tickets sold = s
Adult tickets sold = a
Cost of total tickets sold = 60
Student tickets sold + Adult tickets sold = Total Tickets sold
s+a = 20
s = 20-a .......(1)
Cost of Adult Ticket = 4.5
Cost of Students Ticket = 2
Cost of Student tickets sold = s*2=2s
Cost of Adult tickets sold = a*4.5 =4.5a
Cost of Student tickets sold + Cost Adult tickets sold = Cost of Total Tickets sold
2s+4.5a=60
Put the value of s in above equation
2(20-a)+4.5a=60
40-2a+4.5a=60
40+2.5a=60
2.5a=60-40
2.5a=60-40
2.5a=20
Divide by 2.5
2.5a/2.5=20/2.5
a=8
Put the value of a in (1)
s = 20-a .......(1)
s = 20-8
s = 12
Student tickets sold = 12
Adult tickets sold = 8
Check
60=12*2 + 8*4.5
60=24+36
60=60